TOWARDS AN ACCURATE INFRARED LINELIST FOR SO_{2}

\underline{XINCHUAN HUANG}, SETI Institute, 189 Bernardo Ave, Suite #100, Mountain View, CA, 94043; DAVID W. SCHWENKE, MS T27B-1, NASA Ames Research Center, Moffett Field, CA, 94035; TIMOTHY J. LEE, MS 245-1, NASA Ames Research Center, Moffett Field, CA, 94035.

The ”Best Theory + High-resolution Expt Data” strategy now extends from NH_{3}, CO_{2} to SO_{2} which is considered a ”weed” in high resolution astronomical data analysis (such as for Herschel, SOFIA, and JWST). Refining a high quality ab initio potential energy surface (PES) with selected HITRAN data, we have made significant progress toward an accurate IR line list for SO_{2}. Similar to the situation for CO_{2}, we have achieved 0.01-0.02 cm^{−1} accuracy for reliable line positions. Outliers in existing HITRAN models have been identified. Compared to CO_{2}, the SO_{2} rovibrational energy levels are more difficult to converge. A pure, experimentally measured dataset (instead of HITRAN models) would greatly enhance the reliabilty of higher J/K and higher energy levels computed on the refined PES. An initial IR line list has been generated to test an ab initio dipole moment surface (DMS). Limitations, deficiencies, and future developments for both line positions and IR intensities will be discussed.